# Wave Function Collapse, Lorentz Invariance, and the Third Postulate of Relativity

### v1

@@ Abstract
In this work it is shown that this modified Schrödinger equation is naturally Lorentz invariant. It is further argued that the additional spacetime structures that it requires provide a way to implement the assumption that spacelike-separated operators (and measurements) commute, and that this assumption of local commutativity should be regarded as a third postulate of relativity.
@@ I Introduction
Gisin’s work was one of a number of proposed stochastic modifications of the Schrödinger equation aimed at resolving the measurement problem[2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15]. Most of these proposals are designed to collapse the state vector to either an approximate position state or to an energy eigenstate. These attempted solutions have often been met with skepticism because they introduce new, ad hoc physical constants and imply small violations of conservation laws. A recent work has shown how to eliminate these problematic features[16]. It is based on the idea that wave function collapse is induced by the elementary interactions that establish correlations between physical systems, and was motivated by the fact that these correlating interactions play a central role both in the measurement of physical quantities and in the instantiation and transmission of physical information. In this work it will also be shown that this proposed equation maintains Lorentz invariance in a natural way.
@@
Concerns about nonlocality were first raised almost immediately after the Schrödinger equation was proposed (by Einstein at the 1927 Solvay conference. See also [17].) However, they were really sharpened by Bell when he showed that the correlations between entangled systems that are separated by spacelike intervals could not be explained by any account in which all physical processes are restricted to propagate only within the light cone[18]. Bell’s argument relied on entanglement between two elementary systems. Realistic measurement situations involve entanglement among many more systems, some of which are macroscopic. This large scale entanglement magnifies the effect of the nonlocality. This is due to the fact that measurement collapses the system to an entire entangled branch of the wave function, and not simply to an eigenstate of the measured system.
@@
As suggested by Weinberg’s very careful characterization of local commutativity described above, the kind of conflation demonstrated in the quotation is simply misguided. The assumption of the commutativity of spacelike-separated operators entails substantially more than a limitation on the speed of light. It implies that any physical effects across spacelike intervals must not transmit physical information. [^1] The use of the notion of “physical information” in formulating the no-superluminal signaling principle is based on the characterization of information as a reproducible and referential property of physical systems given in [25]. This is the reason that the assumption is made. It involves an implicit recognition that there are nonlocal effects, and that there is the need to regulate them. But, this kind of regulation ought to be explained, and not simply imposed by fiat.
As mentioned above Einstein’s postulate about the invariance of the speed of light is implemented by attributing a light cone structure to spacetime. In other words, this postulate is explained as a consequence of the fundamental nature of spacetime. Should we not then also consider adding structure to spacetime to explain how nonlocal effects propagate and how they are regulated? This is exactly what stochastic collapse equations do. By assuming a foliation of spacetime into spacelike surfaces, and invoking a stochastic process (or processes) they provide the desired explanation. By making wave function collapse and the Born rule follow from the fundamental equation of the theory, they provide a coherent logical structure for contemporary physics.
@@
The next section describes the basic assumptions employed by nonlinear, norm-preserving stochastic collapse equations and illustrates how they work. Section III reviews the proposal of [16], which eliminates the need to introduce new ad hoc physical constants and insures that conservation laws are respected in individual experiments. It then goes on to demonstrate that it meets a critical test of Lorentz invariance in a very natural way. Section IV addresses the question of what constitutes “serious” or natural Lorentz invariance.
@@ II Stochastic Collapse Equations
This section first provides an explanation of why stochastic collapse equations need to add structures to spacetime, specifically a preferred foliation associated with a stochastic process (or processes). This is followed by an illustration of how these dynamic equations generate collapse in accord with the Born rule. The literature cited earlier contains more general and formal demonstrations of how these equations work; the purpose here is just to provide an intuitive understanding based on a simplified case that captures the essential features. The idea is to show that it is not necessary to paste the measurement postulates onto quantum theory in an ad hoc manner, but rather that they follow in a very natural way from a relatively simple modification of the Schrödinger equation that takes into account the fundamentally probabilistic nature of quantum theory as we currently understand it.
@@ II.1 The Need for Additional Spacetime Structure
Interactions play a critical role in stochastic collapse equations. In the proposal of [16] that will be described in detail in Section III they play the central role because it is assumed that it is interactions that actually induce the collapse. In other proposals they are crucial to establishing the large scale entanglement that allows nonlinear collapse to occur on macroscopic scales while leaving the (almost) linear quantum behavior of microscopic systems essentially undisturbed. Entanglement is a generic result of interaction, as shown by Gemmer and Mahler and by Durt[26, 27]. The fact that the changes that occur in measurements involve entire entangled branches of the wave function implies that any proposal that treats wave function collapse as a real physical process and seeks to explain it by relating it to other fundamental processes must add structure to relativistic spacetime.
To see this we need to contrast the idea of wave function collapse as a process with the conventional view that it is an instantaneous event that just happens at some point during (or immediately after) a measurement. If there are two or more spacelike-separated measurements involving some of the same systems the final outcome is independent of the order in which the instantaneous collapse is assumed to occur. This has to be the case in order to insure compatibility with relativity. On the other hand, stochastic collapse equations assume that the collapse process is mediated by entanglement relations, and, therefore, it is essential that the entangled branches be well defined at each stage of the process. Thus, it is necessary that there be some means to determine whether other spacelike-separated systems that have interacted with other branches of the measured wave function are involved in the collapse. A foliation of spacetime into spacelike surfaces provides the sequencing that determines the structure of a wave function when collapse occurs.
The need for a method to sequence spacelike-separated interactions also arises in derivations of the Born probability rule in collapse processes. Again, we can contrast the conventional view of collapse as an instantaneous occurrence with attempts to describe it as a process. In the conventional view one simply assumes the Born rule (or some principle such as local commutativity that implies it). As will be described below, when collapse is assumed to result from small nonlocal changes that occur in the various branches of the wave function it is essential to the derivation of the rule that the changes are proportional to the relative amplitudes of the branches. Each such change results in a redistribution of amplitudes among the branches. Therefore, the sequence of changes, which can occur at spacelike separation, makes a difference in the result.
For these reasons proposals to describe wave function collapse as a process typically assume that additional spacetime structure such as a foliation is required.
@@ II.2 Nonlinear, Norm-Preserving Stochastic Collapse Equations
The binary character of measurement processes noted above allows us to illustrate the essential operation of stochastic collapse equations using a simple two state system. Consider a system with a wave function $\psi$ and Hamiltonian, $\mathbf{H}$. The system evolves under the action of the Hamiltonian and of a stochastic collapse operator that is constructed from a self-adjoint operator, $\mathbf{O}$, with two eigenstates, $|x\rangle$ and $|y\rangle$, associated with eigenvalues $\mathbf{a}$ and $\mathbf{b}$. The eigenstates of the operator, $\mathbf{O}$, define the collapse basis. The nonlinear stochastic operator is defined as: $\,\mathcal{O}\;\equiv\;k[\mathbf{O}-\langle\,\psi|\mathbf{O}|\psi\,\rangle]\,$, where $\langle\,\psi|\mathbf{O}|\psi\,\rangle$ is the expectation value of the operator, $\mathbf{{O}}$, in the state, $\psi$, and $k$ is a constant that helps to determine the scale of the collapse effects and also insures that $\mathcal{O}$ is dimensionless. For example, if $\mathcal{O}$ is based on the position operator $k$ could determine the range of the collapse effect.
@@ III Interaction-Induced Wave Function Collapse
The degree of correlation between two systems that is generated during an interaction depends on the extent to which the interaction changes the individual state of each system. This, in turn, depends on the strength of the interaction and the resistance of each of the systems to a change of state. The strength is measured by the interaction potential energy. This is assumed to depend on the separation between the two systems and to decrease as the separation increases. For systems $j$ and $k$ it will be indicated as $\mathbf{{V}_{jk}}$. The resistance to change depends on the mass of the systems, $m_{j}$ and $m_{k}$. The effective mass of elementary systems such as electrons in bound states is altered by the binding interactions. So atoms, molecules, and larger complex structures are treated as single systems with a total mass and a net charge (or electric multipole moment).
So the collapse operator is based on the interaction potential energies, $\mathbf{{V}_{jk}}$, and is proportional to the ratio, $\mathbf{{V}_{jk}}/(m_{j}+m_{k})$. As mentioned in Section II the stochastic operator must be dimensionless. To convert the denominator to an energy it is multiplied by the square of the speed of light, $c^{2}$. This is the only nonarbitrary speed, it eases the way to a relativistic generalization, and it eliminates the need to introduce an arbitrary constant, $k$. It is also necessary, as in Section II, to subtract the expectation value, $\langle\,\psi|\mathbf{{V}_{jk}}|\psi\,\rangle$. So the collapse operator associated with the interaction between systems $j$ and $k$ is:
@@
In most proposed collapse equations the rate or frequency parameter, $\gamma$, is chosen in an ad hoc fashion to insure that collapse occurs on an appropriate scale. However, since it is assumed here that collapse effects are induced by the physical processes that establish correlations between systems it is possible to define the frequency parameter in terms of the rate at which the correlations are generated. Since correlations are established through the exchange of conserved quantities, and since these exchanges are associated with variations in the interaction potentials, $\mathbf{{V}_{jk}}$ we can define a rate parameter associated with each interaction in terms of the rate at which these potentials vary:
@@
$\gamma_{jk}\;\equiv\;\frac{\parallel\frac{d}{dt}\langle\,\psi|\mathbf{{V}_{jk}}|\psi\,\rangle\parallel}{\parallel\langle\,\psi|\mathbf{{V}_{jk}}|\psi\,\rangle\parallel_{max}}.$ (6)
The numerator in this expression is obviously well defined at every stage of the interaction. The fact that the denominator is also well defined is based on a couple of observations. First, since the Born probability rule follows from the basic structure of collapse equations the expectation value for the full collapse equation is the same as for the Schrödinger equation associated with it. Second, the maximum of the expectation value over the course of the interaction can be projected by integrating the Schrödinger equation. This can be done because the interaction potentials are functions of position, and hence, time-independent. It is true that both numerator and denominator can vary over the course of the interaction, but the crucial point is that they are well defined at every stage. The fact that both numerator and denominator pick out only the interacting components of the the wave function insures that the rate parameter is independent of the amplitude of these components, and that is crucial for guaranteeing compliance with the Born probability rule. The norm of both numerator and denominator is taken in order to insure the positivity of the expression for $\gamma_{jk}$. [^7] This formulation of the rate parameter is simpler and more general than the expression that was used in [16]. Note also that the expression for the denominator can be integrated in both directions in time.
A critical implication of this formulation for $\gamma_{jk}$ is that when multiplied by $dt$ it integrates to a value of order $1$ over the course of the interaction. This is also true for the expression $\sqrt{\gamma_{jk}}d\xi$. This fact allows one to treat each interaction as a discrete event, and makes it possible to estimate the scale on which collapse occurs and the duration of the collapse process. This kind of analysis shows how the scale and duration depend on the average strength of the interactions involved. For this purpose it is useful to have an estimate of the average interaction rate and duration. This can be done by examining the interaction term in the Schrödinger equation,
$\frac{d\psi}{dt}\,=\,-\,\frac{i\mathbf{{V}_{jk}}}{\hbar}\,\psi.$ (7)
This shows that the average rate is proportional to the interaction strength
$\gamma_{jk}\,\approx\,\frac{\Delta\mathbf{{V}_{jk-max}}}{\hbar}.$ (8)
where $\Delta\mathbf{{V}_{jk-max}}$ is the maximum difference in potential energy over the course of the interaction. The characteristic interaction time is the inverse of this quantity:
$dt_{int}\,\equiv\,\frac{\hbar}{\Delta\mathbf{{V}_{jk-max}}},$ (9)
@@
For two electrons separated by the Bohr radius the duration would be about $2.5*10^{-17}$ seconds. Rates and durations for other interactions can be scaled from this estimate, taking into account the mass and charge of the systems involved.
@@
Finally, the full collapse operator is obtained by summing over the operators for every pair, $j,k$:
${\mathcal{V}}\;\equiv\;\sum_{j<k}{\mathcal{V}}_{jk}.$ (10)
@@
A detailed proof that 11 results in collapse with the correct probabilities is given in [16], along with estimates of the scale and duration of collapse processes. As indicated these depend on the average strength of the interactions involved. Such processes can involve anywhere from about $10^{8}$ to $10^{16}$ elementary interactions. Since many of the interactions can be occurring in parallel the durations are typically very small fractions of a second as can be seen from the estimates of the durations associated with individual interactions given above.
@@
This review of the proposal in [16] is intended as background for the demonstration that 11 is Lorentz invariant. The general question of Lorentz invariance hangs over any proposal that deals with the measurement problem. But there is a particular concern that arises in connection with the type of collapse equation discussed here because they are formulated in a preferred reference frame. There is essentially zero probability that any reference frame in which one chooses to analyze a collapse process will coincide with the preferred frame. The critical question is whether the magnitude of the collapse effects is independent of the frame in which they are considered.
@@
IV “Serious” Lorentz Invariance
What is meant by “serious” Lorentz invariance? One common attitude is that no proposal that requires adding structure to spacetime beyond a Lorentzian metric can be considered to be genuinely relativistic. It implies that any new physics must work within the limited framework that was erected in a classical context prior to the development of quantum theory.

### v2

@@ Abstract
In this work it is shown that this modified Schrödinger equation is Lorentz invariant. It is further argued that the additional spacetime structure that it requires provides a way to implement the assumption that spacelike-separated operators (and measurements) commute, and that this assumption of local commutativity should be regarded as a third postulate of relativity.
@@ I Introduction
Gisin’s work was one of a number of proposed stochastic modifications of the Schrödinger equation aimed at resolving the measurement problem[2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15]. Most of these proposals are designed to collapse the state vector to either an approximate position state or to an energy eigenstate. These attempted solutions have often been met with skepticism because they introduce new, ad hoc physical constants and imply small violations of conservation laws. A recent work has shown how to eliminate these problematic features[16]. It is based on the idea that wave function collapse is induced by the elementary interactions that establish correlations between physical systems, and was motivated by the fact that these correlating interactions play a central role both in the measurement of physical quantities and in the instantiation and transmission of physical information. In this work it will also be shown that this proposed equation maintains Lorentz invariance.
@@
Concerns about nonlocality were first raised almost immediately after the Schrödinger equation was proposed (by Einstein at the 1927 Solvay conference; see also [17].) However, they were really sharpened by Bell when he showed that the correlations between entangled systems that are separated by spacelike intervals could not be explained by any account in which all physical processes are restricted to propagate only within the light cone[18]. At least in some sense, measurements do affect spacelike-separated systems.
@@
As suggested by Weinberg’s very careful characterization of local commutativity described above, the kind of conflation demonstrated in the quotation is simply wrong. The assumption that spacelike-separated operators commute is made to insure that any effects on physical systems that do propagate across spacelike intervals do not transmit any physical information. It involves the (at least) implicit recognition that there are nonlocal effects, and that there is a need to regulate them. But, this kind of regulation ought to be explained, and not simply imposed by fiat.
As mentioned above Einstein’s postulate about the invariance of the speed of light is implemented by attributing a light cone structure to spacetime. In other words, this postulate is explained as a consequence of the fundamental nature of spacetime. Should we not then also consider adding structure to spacetime to explain how nonlocal effects propagate and how they can be regulated to maintain Lorentz invariance? This is exactly what stochastic collapse equations do. By assuming a foliation of spacetime into spacelike surfaces, and invoking a stochastic process (or processes) they provide the desired explanation. By making wave function collapse and the Born rule follow from the fundamental equation of the theory, they provide a coherent logical structure for contemporary physics.
@@
The next section describes the basic assumptions employed by nonlinear, norm-preserving stochastic collapse equations and illustrates how they work. Section III reviews the proposal of [16], which eliminates the need to introduce new ad hoc physical constants and insures that conservation laws are respected in individual experiments. It then goes on to demonstrate that it meets a critical test of Lorentz invariance in a very natural way. Section IV addresses the question of what it means for a theory to be ”genuinely relativistic.”
@@ II Stochastic Collapse Equations
This section first provides an explanation of why stochastic collapse equations need to add structures to spacetime, specifically a preferred foliation associated with a stochastic process (or processes). This is followed by an illustration of how these dynamic equations generate collapse in accord with the Born rule. The literature cited earlier contains more general and formal demonstrations of how these equations work; a proof that the equation described in Section III entails collapse in conformity to the Born rule has been presented in [16]. The purpose here is just to provide an intuitive understanding based on a simplified case that captures the essential features. The idea is to show that it is not necessary to paste the measurement postulates onto quantum theory in an ad hoc manner, but rather that they follow in a very natural way from a relatively simple modification of the Schrödinger equation that takes into account the fundamentally probabilistic nature of quantum theory as we currently understand it.
@@ II.1 The Need for Additional Spacetime Structure
Interactions play a critical role in stochastic collapse equations. In the proposal of [16] that will be described in detail in Section III they play the central role because it is assumed that it is interactions that actually induce the collapse. In other proposals they are crucial to establishing the large scale entanglement that allows nonlinear collapse to occur on macroscopic scales while leaving the (almost) linear quantum behavior of microscopic systems essentially undisturbed. Because these equations typically require an extremely large number of elementary systems to become entangled in order to generate collapse it is essential that the entanglement relations are well defined throughout the collapse process. To insure this there must be some means of sequencing the interactions that generate these relations. In typical measurement processes some of these interactions are spacelike-separated. The assumption of a preferred reference frame (or more generally, a foliation of spacetime) provides the necessary sequencing.
@@ II.2 Nonlinear, Norm-Preserving Stochastic Collapse Equations
The binary character of measurement processes noted above allows us to illustrate the essential operation of stochastic collapse equations using a simple two state system. Consider a system with a wave function $\psi$ and Hamiltonian, $\mathbf{H}$. The system evolves under the action of the Hamiltonian and of a stochastic collapse operator that is constructed from a self-adjoint operator, $\mathbf{O}$, with two eigenstates, $|x\rangle$ and $|y\rangle$, associated with eigenvalues $\mathbf{a}$ and $\mathbf{b}$. The eigenstates of the operator, $\mathbf{O}$, define the collapse basis. The nonlinear stochastic operator is defined as: $\,\mathcal{O}\;\equiv\;k[\mathbf{O}-\langle\,\psi|\mathbf{O}|\psi\,\rangle]\,$, where $\langle\,\psi|\mathbf{O}|\psi\,\rangle$ is the expectation value of the operator, $\mathbf{{O}}$, in the state, $\psi$, and $k$ is a constant that helps to determine the strength and scale of the collapse effects and also insures that $\mathcal{O}$ is dimensionless. For example, if $\mathcal{O}$ is based on the position operator $k$ could determine the range of the collapse effect.
@@
$\mathbf{O}\psi\,=\,\mathbf{a}\alpha|x\rangle+\mathbf{b}\beta|y\rangle$, and its expectation value is $\langle\,\psi|\mathbf{O}|\psi\,\rangle\,=\,\mathbf{a}\alpha^{2}\,+\,\mathbf{b}\beta^{2}$. The action of the stochastic operator on the wave function can be expanded as:
@@
I have assumed here that the terms, $k$ and $\gamma$, are constant, and, hence, independent of the amplitudes, $\alpha$ and $\beta$. This independence from the amplitudes is crucial for the derivation of the Born rule. It is this feature that must be maintained when $k$ and $\gamma$ are replaced by parameters that vary in a manner that is dependent on the interaction strength. These parameters are discussed in the next section.
@@ III Interaction-Induced Wave Function Collapse
The degree of correlation between two systems that is generated during an interaction depends on the extent to which the interaction changes the individual state of each system. This, in turn, depends on the strength of the interaction and the resistance of each of the systems to a change of state. The strength is measured by the interaction potential energy. This is assumed to depend on the separation between the two systems and to decrease as the separation, $\mathbf{r}$, increases. For systems $j$ and $k$ it will be indicated as $\mathbf{{V}_{jk}}\,=\,\mathbf{{V}_{jk}}(\mathbf{{r}_{jk}})$. The resistance to change depends on the mass of the systems, $m_{j}$ and $m_{k}$. The effective mass of elementary systems such as electrons in bound states is altered by the binding interactions. So atoms, molecules, and other complex structures are treated as single systems with a total mass and a net charge (or electric multipole moment).
So the collapse operator is based on the interaction potential energies, $\mathbf{{V}_{jk}}$, and is proportional to the ratio, $\mathbf{{V}_{jk}}/(m_{j}+m_{k})$. As mentioned in Section II the stochastic operator must be dimensionless. To convert the denominator to an energy it is multiplied by the square of the speed of light, $c^{2}$. This is the only nonarbitrary speed, it is crucial for maintaining Lorentz invariance, and it eliminates the need to introduce an arbitrary constant, $k$. It is also necessary, as in Section II, to subtract the expectation value, $\langle\,\psi|\mathbf{{V}_{jk}}|\psi\,\rangle$. So the component of the collapse operator associated with the interaction between systems $j$ and $k$ is:
@@
In most proposed collapse equations the rate or frequency parameter, $\gamma$, is a constant. It is chosen in a rather ad hoc manner to minimize deviations from linearity at an elementary level while insuring collapse on a macroscopic scale. In contrast, since it is assumed here that collapse effects are induced by the physical processes that establish correlations between systems it is possible to define this parameter in terms of the rate at which the correlations are generated. Since correlations are established through the exchange of conserved quantities, and since these exchanges are associated with variations in the interaction potentials, $\mathbf{{V}_{jk}}$, we can define a rate parameter, $\gamma_{jk}$, associated with each interaction in terms of the rate at which the interaction proceeds. As described in Section II the term, $\sqrt{\gamma_{jk}}$, is multiplied by the stochastic differential, $d\xi$.
The rate can be defined so that it integrates to a value of approximately, $1$, over the course of the interaction, and goes to zero when the interacting systems either settle into a stationary state or separate to a distance at which the interaction effectively ends. This insures that the stochastic collapse equation reduces to the ordinary Schrödinger equation in these situations. This can be done by taking the ratio of the rate of change of potential energy to the maximum magnitude of potential energy that occurs during the interaction:
(rate of change of P.E. / max P.E).
Both the rate of change and the maximum value are affected by the action of the Hamiltonian and the stochastic operator. However, because the stochastic action is unbiased and because the changes that it induces in the numerator and denominator are correlated its effects tend to cancel out. Because collapse equations involve stochastic differentials, $d\xi$, time derivatives are not well defined. But, it is possible to use the Hamiltonian to express its contribution to the rate of change of potential energy in terms of spatial derivatives:
$\begin{array}[]{ll}\Big{|}\Big{|}\;(i\hbar)\,\mathbf{\int}\,\mathbf{{V}_{jk}}\,\Big{(}\,\frac{\psi^{*}\,\mathbf{\nabla_{j}}^{2}\psi\,-\,\psi\mathbf{\nabla_{j}}^{2}\psi^{*}}{2m_{j}}\,+\,\frac{\psi^{*}\,\mathbf{\nabla_{k}}^{2}\psi\,-\,\psi\mathbf{\nabla_{k}}^{2}\psi^{*}}{2m_{k}}\,\Big{)}\;\Big{|}\Big{|}.\end{array}$ (6)
The norm is taken to insure a positive rate.
To establish that the expression for the denominator in $\gamma_{jk}$, $||\langle\,\psi|\mathbf{{V}_{jk}}|\,\psi\rangle||_{max}$, is well defined note that the collapse equation can be integrated both backward and forward in time. Because the Hamiltonian is independent of time it is clear that the associated Scrödinger equation is integrable. The fact that the stochastic operator is unbiased implies that averaging over all of the unravellings of the collapse equation will reproduce the value obtained by the Scrödinger equation integral. Observe also that the stochastic changes in the denominator are matched by those in the numerator; so their overall effect on $\gamma_{jk}$ is to leave it essentially unchanged. It is, of course, not necessary to actually carry out the integration since $\gamma_{jk}$ is designed to integrate to a value of order, $1$, over the course of the interaction. With this understanding the variable rate parameter is defined as follows:
@@
$\gamma_{jk}\,\equiv\,\frac{||(i\hbar)\,\mathbf{\int}\,\mathbf{{V}_{jk}}\Big{(}\frac{\psi^{*}\mathbf{\nabla_{j}}^{2}\psi-\psi\mathbf{\nabla_{j}}^{2}\psi^{*}}{2m_{j}}+\frac{\psi^{*}\mathbf{\nabla_{k}}^{2}\psi-\psi\mathbf{\nabla_{k}}^{2}\psi^{*}}{2m_{k}}\,\Big{)}||}{||\langle\,\psi|\mathbf{{V}_{jk}}|\,\psi\rangle||_{max}}.$ (7)
An important point to note about 7 is that both numerator and denominator pick out only the interacting components of the the wave function. This insures that $\gamma_{jk}$ is dependent only on the portion of the wave function that is involved in the interaction. This insures that the rate parameter is independent of the amplitudes of the interacting and noninteracting components. This feature is necessary to insure compliance with the Born probability rule as illustrated in the discussion in Section II.
The period over which $\gamma_{jk}$ is significantly different from zero depends on the initial conditions of the interaction. However, the time during which the vast majority of the momentum and energy are exchanged (and the amplitude is transferred) depends essentially on the maximum interaction energy:
$dt_{int}\,\approx\,\frac{\hbar}{\mathbf{{V}_{jk-max}}},$ (8)
@@
For two electrons with a maximum interaction energy equal to the potential at the Bohr radius the duration would be about $2.5*10^{-17}$ seconds. Rates and durations for other interactions can be scaled from this estimate, taking into account the mass and charge of the systems involved.
As promised a few paragraphs back this formulation for $\gamma_{jk}$ implies that when multiplied by $dt$ it integrates to a value of order $1$ over the course of the interaction. This is also true for the expression $\sqrt{\gamma_{jk}}d\xi$. This fact allows one to treat each interaction as a discrete event, and makes it possible to estimate the scale on which collapse occurs and the duration of the collapse process. This kind of analysis shows how the scale and duration depend on the average strength of the interactions involved. The estimates of the average interaction rate and duration were given above.
@@
Finally, the full collapse operator is obtained by multiplying the operators by the square root of the rate parameters and summing over the terms for each $(j,k)$ pair:
${\mathcal{V}}\;\equiv\;\sum_{j<k}{\mathcal{V}}_{jk}\sqrt{\gamma_{jk}}.$ (9)
@@
A detailed proof that 10 results in collapse with the correct probabilities is given in [16], along with estimates of the scale and duration of collapse processes. As indicated these depend on the average strength of the interactions involved. Such processes can involve anywhere from about $10^{8}$ to $10^{16}$ elementary interactions. Since many of the interactions can be occurring in parallel the durations are typically very small fractions of a second as can be seen from the estimates of the durations associated with individual interactions given above.
@@
This review of the proposal in [16] is intended as background for the demonstration that 10 is Lorentz invariant. The general question of Lorentz invariance hangs over any proposal that deals with the measurement problem. But there is a particular concern that arises in connection with the type of collapse equation discussed here because they are formulated in a preferred reference frame. There is essentially zero probability that any reference frame in which one chooses to analyze a collapse process will coincide with the preferred frame. The critical question is whether the magnitude of the collapse effects is independent of the frame in which they are considered.
@@
IV Reconciling Relativity and Quantum Nonlocality
One common attitude is that no proposal that requires adding structure to spacetime beyond a Lorentzian metric can be considered to be genuinely relativistic. It implies that any new physics must work within the limited framework that was erected in a classical context prior to the development of quantum theory.